OpenStudy (anonymous):

What is the value of 4a3 – 2b2 + 5 when a = 3 and b = −2?

OpenStudy (anonymous):

Plug in those numbers. Be sure to follow the order of operations, darling!

OpenStudy (anonymous):

ok i got 105

OpenStudy (anonymous):

@AnnalieseBronz

OpenStudy (anonymous):

=4*27-2*4+5 =108-8+5 =100+5

jabez177 (jabez177):

105 is your answer.